Baird Camps, Inc. leases the land on which it builds camp sites. Baird is considering opening a new site on land that requires $3,400 of rental payment per month. The variable cost of providing service is expected to be $6 per camper. The following chart shows the number of campers Baird expects for the first year of operation of the new site:

Jan. Feb. Mar. Apr. May June July Aug. Sept. Oct. Nov. Dec. Total
410 390 400 420 760 660 800 810 500 530 650 470 6,800


Required

Assuming that Baird wants to earn $10 per camper, determine the price it should charge for a camp site in February and August. (Do not round intermediate calculations.)
Answer #1

Total amount required in February = Fixed rental + Number of campers * Variable cost + Profit per camper* number of camper

= 3400 + 390 * 6 + 390 *10

= 9640

Amount to be charged per camper = 9640 / 390 = $24.72

Total amount required in August = Fixed rental + Number of campers * Variable cost + Profit per camper* number of camper

= 3400 + 810 * 6 + 810 *10

= 16360

Amount to be charged per camper = 16360/810 = $20.20